Each set of ordered pairs represents a function. Write a rule that represents the function.
1. (2,10), (4,20), (5,25), (7,35), (9,45)
2. (2,5), (4,9), (5,11), (7,15), (10,21)
The answer I got the the first set of ordered pairs was f(x) = 5x, and for the second pair of ordered pairs, the answer I got was f(x) = 2x+1. Are my answers correct?

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52797)

.

They are correct, but they are not answers, obtained by you.

They are answers, obtained by other people (by the tutors) for you.

I write it for sure, because it was ME, who obtained these answers #1 for you.

===============

Hello, with your posts, you REALLY surprise me.


Not only you do not understand NEITHER the problem, NOR the solution, NOR the answer;
in addition, you ask about everything again.


It is just far beyond of my understanding.

It means, that after obtaining the solutions and the answers, you do understand NOTHING,
exactly as you did understand nothing before obtaining it !


For what reason do you then write to this forum ?



Answer by math_tutor2020(3817)


I'll show you how to verify one of the points for each function.

=================================================

Problem 1
For the point (2,10), it means x = 2 leads to y = f(x) = 10.

So,
y = 5x
y = 5*2
y = 10
which confirms (2,10) is on the equation y = 5x
I'll let you check the other points for problem 1.

=================================================

Problem 2

Same idea as the previous problem
Let's see if x = 2 leads to y = 5
y = 2x+1
y = 2*2+1
y = 4+1
y = 5
Sure enough it does, so we have confirmed that (2,5) is on the function y = 2x+1
I'll let you check the other points.

You can use graphing tools like Desmos or GeoGebra to visually confirm the answers quickly.
However, I recommend getting practice doing this by hand before you turn to a calculator.